# A note for double Hölder regularity of the hydrodynamic pressure for weak solutions of Euler equations

### v2

@@ Abstract.
We give an elementary proof for the double Hölder regularity of the hydrodynamic pressure for weak solutions of the Euler Equations in a bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d}$; $d\geq 3$. That is, for velocity $u\in C^{0,\gamma}(\Omega;\mathbb{R}^{d})$ with some $20<\gamma<1/$ 2, we show that the pressure $p\in C^{0,2\gamma}(\Omega)$. This is motivated by the studies of turbulence and anolalous dissipation in mathematical hydrodynamics and, recently, has been established in [L. De Rosa, M. Latocca, and G. Stefani, Int. Math. Res. Not. 2024.3 (2024), 2511–2560] over $C^{2,\alpha}$-domains by means of pseudodifferential calculus. Our approach involves only standard elliptic PDE techniques, and relies crucially on a variant of the modified pressure introduced in [C. W. Bardos, D. W. Boutros, and E. S. Titi, Hölder regularity of the pressure for weak solutions of the 3D Euler equations in bounded domains, arXiv: 2304.01952] and the potential estimates in [L. Silvestre, unpublished notes]. The key novel ingredient of our proof is the introduction of two cutoff functions whose localisation parameters are carefully chosen as a power of the distance to ${\partial\Omega}$.
@@ 1. Introduction
Here, $u:[0,T]\times\Omega\to\mathbb{R}^{d}$ is the velocity and $p:[0,T]\times\Omega\to\mathbb{R}$ is the hydrodynamic pressure of the flow, and $\nu:{\partial\Omega}\to\mathbb{R}^{d}$ is the outward normal vector field to the boundary ${\partial\Omega}$. On ${\partial\Omega}$ the impermeability boundary condition is imposed. The goal of this note is to give an alternative, elementary proof for the double Hölder regularity of $p$.
@@ 1.1. Main Theorem
Let $(u,p)$ be a weak solution of the Euler Equation (1.1) in the bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d}$ for $d\geq 3$. Assume that $u$ is of Hölder regularity $C^{0,\gamma}(\Omega;\mathbb{R}^{d})$ in the spatial variable for some $20<\gamma<1/$ 2. Then the hydrodynamic pressure $p$ is of double Hölder regularity $C^{0,2\gamma}(\Omega)$ in the spatial variable. Moreover, for any $x_{1},x_{2}\in\Omega$ it holds that
$|p(x_{1})-p(x_{2})|\leq C|x_{1}-x_{2}|^{2\gamma},$ (1.2)
where the constant $C$ depends only on $d$, $\gamma$, $\|u\|_{C^{0,\gamma}(\Omega)}$, $\Omega$, and $\min\limits_{i=1,2}{\rm dist}(x_{i},{\partial\Omega})$.
Throughout this note, the time dependence of $u$ and $p$ is immaterial, since the PDE satisfied by the pressure is elliptic and has no explicit dependence on time; see (2.8). Here and hereafter, we suppress the time variable $t$ by writing $u(t,x)\equiv u(x)$, $p(t,x)\equiv p(x)$, and the like. Theorem 1.1 should be understood in the kinematic sense; that is, the inequality (1.2) holds for every $t\in[0,T]$.
@@
The constant $C$ in the estimate (1.2) can be chosen more precisely as follows: denoting $\kappa=\min\limits_{i=1,2}{\rm dist}(x_{i},{\partial\Omega})$, one has
$C=C(d,\gamma)C_{\rm geom}\|u\|_{C^{0,\gamma}(\Omega)}^{2}\cdot\begin{cases}1\qquad\qquad\text{ as~{} }|x_{1}-x_{2}|\leq c_{\rm geom}\kappa,\\ \kappa^{-1+2\gamma}\qquad\qquad\text{ otherwise}.\end{cases}$
Here $C_{\rm geom}$ and $c_{\rm geom}$ are two geometrical constants, depending only on the $C^{0}$-norm of the second fundamental form of ${\partial\Omega}$, the intrinsic diameter of $\Omega$ (the infimum of the length of $C^{1}$-curves connecting any two points inside $\Omega$), and the injectivity radius of $\Omega$. Also, in the limiting process $\kappa\searrow 0$,
  • $C$ is “locally uniformly” in $\kappa$ in the sense that for $|x_{1}-x_{2}|\leq c_{\rm geom}\kappa$, the constant $C$ can be chosen independently of $\kappa$; and
  • for general $x_{1}$, $x_{2}$, the constant may blow up at a rate of $\mathcal{O}\left(\kappa^{-1+2\gamma}\right)$ when approaching the boundary.
@@
Bardos and Titi initiated in [3] the project of extending the Onsager conjecture and related results to bounded domains $\Omega\subset\mathbb{R}^{d}$. The Hölder regularity of the hydrodynamic pressure $p$ in the Euler Equations (1.1) plays a fundamental role in the study of anomalous dissipation on bounded domains [3, 5, 19, 20]. Assuming $u\in C^{0,\gamma}(\Omega)$ with $10<\gamma<$ 1 and $\Omega$ is a bounded $C^{3}$-domain, Bardos and Titi obtained that $p\in C^{0,\gamma}(\Omega)$ for $\Omega\subset\mathbb{R}^{2}$ in [4] and, together with Boutros, proved the same result for $\Omega\subset\mathbb{R}^{3}$ in [2]. Then, utilising among others the techniques developed in [7, 8], De Rosa, Latocca and Stefani [13] established the almost double Hölder regularity in the sense that $p\in C^{1,\min(\alpha,2\gamma-1)}(\Omega)$ for $11/2<\gamma<$ 1 on $C^{2,\alpha}$-domains, and $p\in C^{2\gamma-\varepsilon}(\Omega)$ with arbitrarily small $\varepsilon>0$ for $20<\gamma<1/$ 2 on $C^{3,\alpha}$-domains. The same authors later in [14] extended the double Hölder regularity $p\in C^{2\gamma}(\Omega)$ to the index range $20<\gamma<1/$ 2 on $C^{2,1}$-domains. In addition, for the critical index $\gamma=1/2$, it is proved in [14, Proposition 2.5] that $p\in C^{1}_{\star}(\Omega)$, the Calderón–Zygmund space. This improves the log-Lipschitz regularity of $p$ obtained earlier in Constantin [9].
The approach in the recent work [14] by De Rosa, Latocca and Stefani involves a delicate use of the pseudodifferential calculus and the Littlewood–Paley theory. In this note, on a bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d\geq 3}$, we recover the main result in [14], i.e., the double Hölder regularity $p\in C^{2\gamma}(\Omega)$ for $20<\gamma<1/$ 2, utilising only elementary techniques of (Neumann) Green functions and integration by parts.
Our arguments make use of ideas from Silvestre [22], Bardos, Boutros and Titi [2], and De Rosa, Latocca and Stefani [13, 14]. As in [2] (with an extra correction term), we introduce a modified pressure Weierstrass-$\wp$ (see §2 for details), which equals
@@
Rather than working directly with Weierstrass-$\wp$, we introduce a second cutoff function ${\eta^{\delta}}$ (see Equation (3.4)) which equals 0 in the $\delta/2$-collar of ${\partial\Omega}$ and equals 1 outside the $\delta$-collar. Thus, by working with the elliptic PDE for ${\eta^{\delta}}\wp$ — which is localised to the interior of $\Omega$ — we circumvent the issues caused by highly irregular boundary values (see Remark 2.1 and [2, Section 8]). The cost to pay is that the singular terms $|\nabla{\eta^{\delta}}|\approx\delta^{-1}$ and $|D^{2}{\eta^{\delta}}|\approx\delta^{-2}$ enter our estimates. We overcome this issue by carefully adjusting the “boundary layers” induced by cutoff functions. For reasons that will become transparent along the proof, when estimating $|\wp(x_{1})-\wp(x_{2})|$ for $x_{1}$, $x_{2}$ sufficiently close to each other and $|x_{1}-x_{2}|\approx{\rm dist}\left(\frac{x_{1}+x_{2}}{2},{\partial\Omega}\right)$, we choose
@@
With the above preparations, an adaptation of the representation formula for $p$ in Silvestre [22] (with the integral kernel therein ${\psi_{x_{1},x_{2}}}$ replaced by ${\eta^{\delta}}{\psi_{x_{1},x_{2}}}$, where ${\psi_{x_{1},x_{2}}}$ is the difference between the Neumann Green functions on $\Omega$ with singularities at $x_{1}$ and $x_{2}$) allows us to obtain the double Hölder regularity Theorem 1.1 via direct potential estimates. It is crucial to our arguments that the previous choice of cutoff functions ${\varphi^{\delta}},{\eta^{\delta}}$ ensures that the constant in the inequality (1.2) is locally uniform in $\delta$ up to the boundary. See Remark 2.1 for details.
@@ 1.2. Notations
A continuous function $u:\Omega\to\mathbb{R}$ is $\alpha$-Hölder ($10<\alpha<$ 1) if the following seminorm is finite:
$[u]_{C^{0,\alpha}(\Omega)}:=\sup_{x\neq y\,\text{in }\Omega}\left\{\frac{|u(x)-u(y)|}{|x-y|^{\alpha}}\right\}.$
@@
$\|u\|_{C^{0,\alpha}(\Omega)}:=[u]_{C^{0,\alpha}(\Omega)}+\|u\|_{C^{0}(\Omega)}.$
@@
${\rm II}=\nabla\nu:\Gamma(T({\partial\Omega}))\times\Gamma(T({\partial\Omega}))\longrightarrow\mathbb{R},$
@@
Einstein’s summation convention is assumed throughout. That is, repeated upper and lower indices are always understood as being summed over. Also, for a constant $C>0$, we write $C=C(a_{1},a_{2},\ldots,a_{n})$ to emphasise that $C$ depends only on the parameters $a_{1},a_{2},\ldots,a_{n}$.
@@ 1.3. Organisation
The main result, Theorem 1.1 will be proved in §3. The novel tool for our proof, namely that the second cutoff function ${\eta^{\delta}}$, is elaborated in §3.1. Estimates for Weierstrass-$\wp$ occupy §§3.2 &\&& 3.3. Finally, the proof of Theorem 1.1 is concluded in §3.4.
@@ 2. The modified pressure
${\varphi^{\delta}}(y):=\phi^{\delta}\left({\rm dist}(y,{\partial\Omega})\right)\qquad\text{for }y\in\Omega.$ (2.1)
@@
where the superscript denotes the projection of a vector field to the tangential direction of ${\partial\Omega}$.
@@
Indeed, the solutions Weierstrass-$\wp$ may differ from the one with average zero by addition of some function whose gradient is a harmonic vector field, which is nontrivial on non-simply-connected $\Omega$. But such a function is in $C^{\infty}(\Omega)\cap C^{2}(\overline{\Omega})$ when $\Omega$ is a $C^{2}$-domain, hence does not affect the double Hölder regularity of Weierstrass-$\wp$.
@@ 3.1. The second cutoff function
See Remark 1.1 for the notion of geometrical constants. Such a choice of $x_{1}$ and $x_{2}$ is possible because $\Omega$, as a bounded $C^{2}$-domain, satisfies the uniform interior sphere condition. That is, there exists a uniform constant $r_{0}>0$ depending only on the $C^{2}$-geometry of $\Omega$ such that for each $x\in{\partial\Omega}$, the open ball of radius $r_{0}$ tangent to ${\partial\Omega}$ at $x$ lies entirely in $\Omega$.
The reason for choosing $\beta:=\frac{d-2}{d-2+2\gamma}$ will become transparent later (see the formula (3.7) in the proof of Lemma 3.1). The further restriction $|x_{1}-x_{2}|\geq c_{\Omega}\delta$ in (3.1) will not be invoked until Step 4 of the proof of Lemma 3.2. This restriction is only temporary and shall be removed in the first step of the proof of Theorem 1.1 in §3.4.
@@
${\eta^{\delta}}(y):=\varpi^{\delta}\left({\rm dist}(y,{\partial\Omega})\right)\qquad\text{for }y\in\Omega.$ (3.4)
@@
where we have used the definition of ${\psi_{x_{1},x_{2}}}$ given in (3.2). But by construction ${\bf supp}\,{\mathcal{Q}}\subset\Omega\setminus\Omega_{2\delta}$, disjoint from $\Omega_{10\cdot\delta^{\beta}}$, which implies that

### v3

@@ Abstract.
We give an elementary proof for the interior double Hölder regularity of the hydrodynamic pressure for weak solutions of the Euler Equations in a bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d}$; $d\geq 3$. That is, for velocity $u\in C^{0,\gamma}(\Omega;\mathbb{R}^{d})$ with some $20<\gamma<1/$ 2, we show that the pressure $p\in C^{0,2\gamma}_{\rm int}(\Omega)$. This is motivated by the studies of turbulence and anomalous dissipation in mathematical hydrodynamics and, recently, has been established in [L. De Rosa, M. Latocca, and G. Stefani, Int. Math. Res. Not. 2024.3 (2024), 2511–2560] over $C^{2,1}$-domains by means of pseudodifferential calculus. Our approach involves only standard elliptic PDE techniques, and relies on a variant of the modified pressure introduced in [C. W. Bardos, D. W. Boutros, and E. S. Titi, Hölder regularity of the pressure for weak solutions of the 3D Euler equations in bounded domains, Arch. Rational Mech. Anal. 249 (2025), 28] and the potential estimates in [L. Silvestre, unpublished notes]. The key novel ingredient of our proof is the introduction of two cutoff functions whose localisation parameters are carefully chosen as a power of the distance to ${\partial\Omega}$.
@@ 1. Introduction
Here, $u:[0,T]\times\Omega\to\mathbb{R}^{d}$ is the velocity and $p:[0,T]\times\Omega\to\mathbb{R}$ is the hydrodynamic pressure of the flow, and $\nu:{\partial\Omega}\to\mathbb{R}^{d}$ is the outward normal vector field to the boundary ${\partial\Omega}$. On ${\partial\Omega}$ the impermeability boundary condition is imposed. The goal of this note is to give an alternative, elementary proof for the interior double Hölder regularity of $p$, and to provide refined characterisations for the limiting behaviour when approaching the boundary.
@@ 1.1. Main Theorem
Let $(u,p)$ be a weak solution of the Euler equation (1.1) in the bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d}$ for $d\geq 3$. Assume that $u$ is of Hölder regularity $C^{0,\gamma}(\Omega;\mathbb{R}^{d})$ in the spatial variable for some $20<\gamma<1/$ 2. Then the hydrodynamic pressure $p$ is locally of double Hölder regularity $C^{0,2\gamma}_{\rm loc}(\Omega)$ in the spatial variable. Moreover, for any subdomain ${\widetilde{\Omega}}\Subset\Omega$ and any $x_{1},x_{2}\in{\widetilde{\Omega}}$,
$|p(x_{1})-p(x_{2})|\leq C|x_{1}-x_{2}|^{2\gamma}$ (1.2)
with the constant $C=C\left(d,\gamma,{\widetilde{\Omega}}\right)\cdot\|u\|_{C^{0,\gamma}(\Omega)}^{2}$, where $C\left(d,\gamma,{\widetilde{\Omega}}\right)$ depends only on $d$, $\gamma$, the $C^{2}$-geometry of $\Omega$, and the distance from ${\widetilde{\Omega}}$ to $\partial\Omega$.
@@
In the above theorem, more precisely, one may replace the “$C^{2}$-geometry of $\Omega$” with the following geometrical data: the $C^{0}$-norm of the second fundamental form of ${\partial\Omega}$, the intrinsic diameter of $\Omega$ (i.e., the infimum of the length of $C^{1}$-curves connecting any two points inside $\Omega$), and the injectivity radius of $\Omega$. If a constant $C_{\rm geom}$ depends on these data only, we shall say that $C_{\rm geom}$ is a geometrical constant.
Remark 1.2.
The constant $C$ in Theorem 1.1, (1.2) may be chosen as follows. Setting $\kappa:=\min\Big{\{}{\rm dist}(x_{1},\partial\Omega),{\rm dist}(x_{2},\partial\Omega)\Big{\}}>0$, we have that
$C\approx\begin{cases}1,\qquad\text{if }|x_{1}-x_{2}|\lesssim\kappa^{1+\frac{2\gamma}{d-2}}\text{ or }|x_{1}-x_{2}|\gtrsim 1;\\ \kappa^{\frac{-2\gamma(d-2+2\gamma)}{d-2}},\qquad\text{if }\kappa^{1+\frac{2\gamma}{d-2}}\lesssim|x_{1}-x_{2}|\lesssim 1.\end{cases}$
All symbols $\approx$, $\lesssim$ here are understood as modulo constants depending only on $d$, $\gamma$, and the $C^{2}$-geometry of $\Omega$.
Throughout this note, we suppress the time variable by writing $u(t,x)\equiv u(x)$, $p(t,x)\equiv p(x)$, and the like. Theorem 1.1 should be understood in the kinematic sense; that is, the inequality (1.2) holds for every $t\in[0,T]$.
@@
Bardos and Titi initiated in [3] the project of extending the Onsager conjecture and related results to bounded domains $\Omega\subset\mathbb{R}^{d}$. The Hölder regularity of the hydrodynamic pressure $p$ in the Euler equations (1.1) plays a fundamental role in the study of anomalous dissipation on bounded domains [3, 5, 19, 20]. Assuming $u\in C^{0,\gamma}(\Omega)$ with $10<\gamma<$ 1 and $\Omega$ is a bounded $C^{3}$-domain, Bardos and Titi obtained that $p\in C^{0,\gamma}_{\rm int}(\Omega)$ for $\Omega\subset\mathbb{R}^{2}$ in [4] and, together with Boutros, proved the same result for $\Omega\subset\mathbb{R}^{3}$ in [2]. Then, utilising among others the techniques developed in [7, 8], De Rosa, Latocca and Stefani [13] established the almost double Hölder regularity in the sense that $p\in C^{1,\min(\alpha,2\gamma-1)}(\Omega)$ for $11/2<\gamma<$ 1 on $C^{2,\alpha}$-domains, and $p\in C^{2\gamma-\varepsilon}(\Omega)$ with arbitrarily small $\varepsilon>0$ for $20<\gamma<1/$ 2 on $C^{3,\alpha}$-domains. The same authors later in [14] extended the double Hölder regularity $p\in C^{2\gamma}(\Omega)$ to the index range $20<\gamma<1/$ 2 on $C^{2,1}$-domains. In addition, for the critical index $\gamma=1/2$, it is proved in [14, Proposition 2.5] that $p\in C^{1}_{\star}(\Omega)$, the Calderón–Zygmund space. This improves the log-Lipschitz regularity of $p$ obtained earlier in Constantin [9]. The interior double Hölder regularity of $p$ on $C^{4}$-domains in $\mathbb{R}^{3}$ has also been proved in [2].
The approach in the recent work [14] by De Rosa, Latocca and Stefani involves a delicate application of pseudodifferential calculus and the Littlewood–Paley theory. Here, on a bounded $C^{2}$-domain $\Omega\subset\mathbb{R}^{d\geq 3}$, we prove the interior double Hölder regularity $p\in C^{2\gamma}_{\rm loc}(\Omega)$ for $20<\gamma<1/$ 2, utilising only elementary techniques of (Neumann) Green functions and integration by parts. Moreover, we give explicit bounds on the blowup rate for $\frac{|p(x)-p(y)|}{|x-y|^{2\gamma}}$ as $x,y$ approach the boundary ${\partial\Omega}$. The issue of double Hölder regularity of $p$ up to the boundary might have remained to be clarified, as indicated in [2, p.8, at the end of § 1.3].
Our arguments make essential use of ideas from Silvestre [22] and Bardos–Boutros–Titi [2]. As in [2] (with an extra correction term), we introduce a modified pressure Weierstrass-$\wp$ (see §2 for details), which equals
@@
Rather than working directly with Weierstrass-$\wp$, we introduce a second cutoff function ${\eta^{\delta}}$ (see (3.4)) which equals zero in the $\delta/2$-collar of ${\partial\Omega}$ and equals one outside the $\delta$-collar. Thus, by working with the elliptic PDE for ${\eta^{\delta}}\wp$ — which is localised to the interior of $\Omega$ — we circumvent the issues caused by highly irregular boundary values (see Remark 2.1 and [2, Section 8]). The cost to pay is that the singular terms $\left|\nabla{\eta^{\delta}}\right|\approx\delta^{-1}$ and $\left|D^{2}{\eta^{\delta}}\right|\approx\delta^{-2}$ enter our estimates. We overcome this issue by carefully adjusting the “boundary layers” induced by cutoff functions. For reasons that will become transparent along the proof, when estimating $|\wp(x_{1})-\wp(x_{2})|$ for $x_{1}$, $x_{2}$ sufficiently close to each other and $|x_{1}-x_{2}|\approx{\rm dist}\left(\frac{x_{1}+x_{2}}{2},{\partial\Omega}\right)$, we choose
@@
With the above preparations, an adaptation of the representation formula for $p$ in Silvestre [22] (with the integral kernel therein ${\psi_{x_{1},x_{2}}}$ replaced by ${\eta^{\delta}}{\psi_{x_{1},x_{2}}}$, where ${\psi_{x_{1},x_{2}}}$ is the difference between the Neumann Green functions on $\Omega$ with singularities at $x_{1}$ and $x_{2}$) allows us to obtain the double Hölder regularity Theorem 1.1 via direct potential estimates. It is crucial to our arguments that the previous choice of cutoff functions ${\varphi^{\delta}},{\eta^{\delta}}$ ensures that the constant in the inequality (1.2) is uniform in $\delta$ up to the boundary. See Remark 2.1 for details.
@@ 1.2. Notations
A continuous function $u:{\widetilde{\Omega}}\to\mathbb{R}^{\ell}$ is $\alpha$-Hölder ($10<\alpha<$ 1) if the following seminorm is finite:
$[u]_{C^{0,\alpha}\left({\widetilde{\Omega}}\right)}:=\sup_{x\neq y\,\text{in }{\widetilde{\Omega}}}\left\{\frac{|u(x)-u(y)|}{|x-y|^{\alpha}}\right\}.$
@@
$\|u\|_{C^{0,\alpha}\left({\widetilde{\Omega}}\right)}:=[u]_{C^{0,\alpha}\left({\widetilde{\Omega}}\right)}+\|u\|_{C^{0}\left({\widetilde{\Omega}}\right)}.$
A function $u:\Omega\to\mathbb{R}^{\ell}$ is in $C^{0,\alpha}_{\rm loc}(\Omega)$ if its restriction on ${\widetilde{\Omega}}$ is in $C^{0,\alpha}\left({\widetilde{\Omega}}\right)$ for any ${\widetilde{\Omega}}\Subset\Omega$.
@@
${\rm II}=\nabla\nu:\Gamma\big{(}T({\partial\Omega})\big{)}\times\Gamma\big{(}T({\partial\Omega})\big{)}\longrightarrow\mathbb{R},$
@@
Einstein’s summation convention is assumed throughout. That is, repeated upper and lower indices are always understood as being summed over. Also, for a constant $C>0$, we write $C=C(a_{1},a_{2},\ldots,a_{n})$ to emphasise that $C$ depends only on the parameters $a_{1},a_{2},\ldots,a_{n}$. If $a_{1}=\Omega$ it means that $C$ depends on the $C^{2}$-geometry of $\Omega$ as well as $a_{2},\ldots,a_{n}$.
@@ 1.3. Organisation
The main result, Theorem 1.1, will be proved in §3. The novel tool for our proof, namely that the second cutoff function ${\eta^{\delta}}$, is elaborated in §3.1. Estimates for Weierstrass-$\wp$ occupy §§3.2 &\&& 3.3. Finally, the proof of Theorem 1.1 is concluded in §3.4.
@@ 2. The modified pressure
${\varphi^{\delta}}(y):=\phi^{\delta}\big{(}{\rm dist}(y,{\partial\Omega})\big{)}\qquad\text{for }y\in\Omega.$ (2.1)
@@
where the superscript denotes the projection of a vector field to the tangential plane of ${\partial\Omega}$.
@@
Indeed, the solutions Weierstrass-$\wp$ may differ from the one with average zero by addition of some function whose gradient is a harmonic vector field, which is nontrivial on non-simply-connected $\Omega$. But such a function is in $C^{\infty}(\Omega)$ when $\Omega$ is a $C^{2}$-domain, hence does not affect the double Hölder regularity of Weierstrass-$\wp$.
@@ 3.1. The second cutoff function
See Theorem 1.1 for the notion of geometrical constants. Such a choice of $x_{1}$ and $x_{2}$ is possible because $\Omega$, as a bounded $C^{2}$-domain, satisfies the uniform interior sphere condition. That is, there exists a uniform constant $r_{0}>0$ depending only on the $C^{2}$-geometry of $\Omega$ such that for each $x\in{\partial\Omega}$, the open ball of radius $r_{0}$ tangent to ${\partial\Omega}$ at $x$ lies entirely in $\Omega$.
The reason for choosing $\beta:=\frac{d-2}{d-2+2\gamma}$ will become transparent later (see the formula (3.8) in the proof of Lemma 3.1). The further restriction $|x_{1}-x_{2}|\geq c_{\Omega}\delta$ in (3.1) will not be invoked until Step 4 of the proof of Lemma 3.2.
@@
${\eta^{\delta}}(y):=\varpi^{\delta}\big{(}{\rm dist}(y,{\partial\Omega})\big{)}\qquad\text{for }y\in\Omega.$ (3.4)
@@
where we have used the definition of ${\psi_{x_{1},x_{2}}}$ given in (3.2). But by construction ${\bf supp}\,({\mathcal{Q}})\subset\Omega\setminus\Omega_{2\delta}$, disjoint from $\Omega_{10\cdot\delta^{\beta}}$, which implies that